what persent of 5.4is 1.2?
what % of 20is 3/5?
Meat totaling 125grms contains 25 grms protein and 6 grams fat.what % protein and % fat?

Answer by checkley71(8403)

1.2/5.4=22.22%
-----------------------------------------
(3/5)/20
3/5*1/20
3/100=3%
--------------------------------------------
25/125=.2=20%
-------------------------------------------
6/125=.048=4.8%